OpenStudy (anonymous):

]Which of the following represents the quantity of x to the second minus nine, all over the quantity of x to the second minus three x in simplified form?

OpenStudy (harsimran_hs4):

(x^2 - 9) / (x^2 - 3x) isn`t this a better way to ask a question?? this can be written as (x-3)(x+3)/ x(x-3) clear?? now can you get to the final answer?

OpenStudy (skullpatrol):

$$\frac{(x^2 - 9)}{(x^2 - 3x)}= \frac{(x-3)(x+3)}{x(x-3)} $$

OpenStudy (anonymous):

Harsimram beat me to it. :-) And just for the record, something-or-other "to the second" is stated as something-or-other "squared". Likewise, anything to the 3rd power is "cubed". To the best of my knowledge, squares and cubes are the only roots/powers which enjoy their own special identity, so all other exponents are simply stated "to the power".

OpenStudy (anonymous):

here are the choices, im still confused about this.. 1 minus 3 over x 1 plus 3 over x x minus 3, all over x x plus 3, all over x

OpenStudy (harsimran_hs4):

cancel the common x-3 factor and you get (x+3)/x = 1 + 3/x clear??

OpenStudy (anonymous):

Look at Skullpatrol's expression.

OpenStudy (anonymous):

is it the 3rd answer?

OpenStudy (harsimran_hs4):

1 and 3 both are correct